Use the Laplace transform to solve the given initial-value problem. y – 3y = 8e2t – 2e-, y(0) = 1, y(0) = -1 y(t) =
Answer #1

given y - 3y = 8227 -2e- 410)=1 4110) = -1 fake laplace toans from [82718) - Y10) - 7/0)] -3 [3Y19) - Y(0)) = 8 L (ezt) . -33-532+88+20 = A(3-3)(8-2)(3+1)+B$(8-2)(4+1) + C (8-3)(3+1)8+D $ (8-3)(3-2) put s=o 03-5X0 + 8x0+20 = AX(-3)(-2)(1) +BX0 + CXTake unverse laplace trains form Y(+)- 104() + 4 4 (5-3)-42(12)-3216541) ly(t) = 10 + 11 e 34 _ yetzt - et) u(4) Answer